Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

envelope in a letter

Expert replies
by maihuna » Fri Aug 14, 2009 3:18 am
Three letters are dictated to three persons and an evelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.

2/9
3/9
4/9
1/3
2/3
Charged up again to beat the beast :)
Join the discussion
Source: — Problem Solving |

Re: envelope in a letter

by dtweah » Fri Aug 14, 2009 4:05 am
maihuna wrote:Three letters are dictated to three persons and an evelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.

2/9
3/9
4/9
1/3
2/3
1-P(no letter in proper envelope)
1- (3C1 x 2)/9=1-2/3=1/3

Choose D
Join the discussion

Re: envelope in a letter

by maihuna » Fri Aug 14, 2009 4:25 am
dtweah wrote:
maihuna wrote:Three letters are dictated to three persons and an evelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.

2/9
3/9
4/9
1/3
2/3
1-P(no letter in proper envelope)
1- (3C1 x 2)/9=1-2/3=1/3

Choose D
Nope Nope Big No, ans is definetly wrong, see there will not be 9 combination, see below
1 st letter have 3 option
2nd letter have 2 option
3rd letter have 1 option

so total event space is only 3*2*1 = 6 not 9 as it will be when enumerating E1,E2,E3 with L1,L2,L3.
Charged up again to beat the beast :)
Join the discussion

Re: envelope in a letter

by dtweah » Fri Aug 14, 2009 4:36 am
maihuna wrote:
dtweah wrote:
maihuna wrote:Three letters are dictated to three persons and an evelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.

2/9
3/9
4/9
1/3
2/3
1-P(no letter in proper envelope)
1- (3C1 x 2)/9=1-2/3=1/3

Choose D
Nope Nope Big No, ans is definetly wrong, see there will not be 9 combination, see below
1 st letter have 3 option
2nd letter have 2 option
3rd letter have 1 option

so total event space is only 3*2*1 = 6 not 9 as it will be when enumerating E1,E2,E3 with L1,L2,L3.
E1 L1
E1 L2
E1 L3
E2 L1
E2 L2
E2 L3
E3 L1
E3 L2
E3 L3
Join the discussion

Re: envelope in a letter

by maihuna » Fri Aug 14, 2009 4:50 am
dtweah wrote:
maihuna wrote:
dtweah wrote:
maihuna wrote:Three letters are dictated to three persons and an evelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.

2/9
3/9
4/9
1/3
2/3
1-P(no letter in proper envelope)
1- (3C1 x 2)/9=1-2/3=1/3

Choose D
Nope Nope Big No, ans is definetly wrong, see there will not be 9 combination, see below
1 st letter have 3 option
2nd letter have 2 option
3rd letter have 1 option

so total event space is only 3*2*1 = 6 not 9 as it will be when enumerating E1,E2,E3 with L1,L2,L3.
E1 L1
E1 L2
E1 L3
E2 L1
E2 L2
E2 L3
E3 L1
E3 L2
E3 L3
DTWEAH,
Thats what I was pointing, an envelope in a letter is different from an E's L's combination, pick three letter and three envelope and experiment, how many arrangements will be possible? First letter has three envelope to go, once you do that you are left with two envelope and two letter, so an letter have 2 option, for last one there is only one option, so combining one gets onbly six option.

Out of these 6 in following way one or more letter may go to correct place:

(E1, L1) (E2, L3) (E3, L2)

similarly other two may be choosen, i.e. E2,L2 and E3 L3 in one way only

so for one letter in correct place there are 3 ways

For two letter, if two letter are in correct place third will be automatically in correct place, so only one way of getting it right for more than one..

so a total of 3+1 = 4 ways of getting it right

Reqd Prob = 4/6 = 2/3
Charged up again to beat the beast :)
Join the discussion

Re: envelope in a letter

by dtweah » Fri Aug 14, 2009 4:58 am
maihuna wrote:
dtweah wrote:
maihuna wrote:
dtweah wrote:
maihuna wrote:Three letters are dictated to three persons and an evelope is addressed to each of them, the letters are inserted into the envelopes at random so that each envelope contains exactly one letter. Find the probability that at least one letter is in its proper envelope.

2/9
3/9
4/9
1/3
2/3
1-P(no letter in proper envelope)
1- (3C1 x 2)/9=1-2/3=1/3

Choose D
Nope Nope Big No, ans is definetly wrong, see there will not be 9 combination, see below
1 st letter have 3 option
2nd letter have 2 option
3rd letter have 1 option

so total event space is only 3*2*1 = 6 not 9 as it will be when enumerating E1,E2,E3 with L1,L2,L3.
E1 L1
E1 L2
E1 L3
E2 L1
E2 L2
E2 L3
E3 L1
E3 L2
E3 L3
DTWEAH,
Thats what I was pointing, an envelope in a letter is different from an E's L's combination, pick three letter and three envelope and experiment, how many arrangements will be possible? First letter has three envelope to go, once you do that you are left with two envelope and two letter, so an letter have 2 option, for last one there is only one option, so combining one gets onbly six option.

Out of these 6 in following way one or more letter may go to correct place:

(E1, L1) (E2, L3) (E3, L2)

similarly other two may be choosen, i.e. E2,L2 and E3 L3 in one way only

so for one letter in correct place there are 3 ways

For two letter, if two letter are in correct place third will be automatically in correct place, so only one way of getting it right for more than one..

so a total of 3+1 = 4 ways of getting it right

Reqd Prob = 4/6 = 2/3
By Fundamental Counting Principle If you have 3 shirts and 3 trousers, the number of outfits you can wear is 9. Three Letters and Three envelopes are no different than 3 trousers and and 3 shirts. By your reasoning if I pick one shirt, I am left with only 2 shirts. Total choices are not calculated that way. Wrong. You have 3 choices of Evelopes and letters for your first choice. If you match evelope to letter put it back and examine the other choices. Why are you leaving letters in envelopes to examine choices? We are interested in ALL possibilites. Do the experiment. You will have 9 choices as outlined.
Join the discussion

by maihuna » Fri Aug 14, 2009 5:09 am
Humm...your answer is incorrect, I am seeing where you are making mistake, but I will not point that out, as your attitude is very un-accepting sort of....see yourself then where your fundamental principle of counting is putting you in trouble....
Charged up again to beat the beast :)
Join the discussion

by dtweah » Fri Aug 14, 2009 5:14 am
maihuna wrote:Humm...your answer is incorrect, see yourself then where your fundamental principle of counting is putting you in trouble....
"An Envelope in a letter is different from an E's L's Combination." Why?

E1 E2 E3 L1 L2 L3

L1 can go in E1 , that is one possibility

L1 can go into E2, that is another possibility. What prevents this in the wording of the problem?

L1 can go into L3, a third possibility. What prevents this?
Join the discussion

by dtweah » Fri Aug 14, 2009 5:20 am
maihuna wrote:Humm...your answer is incorrect, I am seeing where you are making mistake, but I will not point that out, as your attitude is very un-accepting sort of....see yourself then where your fundamental principle of counting is putting you in trouble....
Attitude? What has that got to do with math. The FCP is a Law. I stand by it.
Join the discussion

by maihuna » Fri Aug 14, 2009 6:26 am
dtweah wrote:
maihuna wrote:Humm...your answer is incorrect, I am seeing where you are making mistake, but I will not point that out, as your attitude is very un-accepting sort of....see yourself then where your fundamental principle of counting is putting you in trouble....
Attitude? What has that got to do with math. The FCP is a Law. I stand by it.
you mean answer is incorrect? what does it mean by your standing or sleeping or walking with FCP? Does it have any significance...?
Charged up again to beat the beast :)
Join the discussion

by ankitns » Fri Aug 14, 2009 8:19 am
You are correct about the 3 shirt - 3 pants combinations...there will be 3*3 = 9 combinations....but this problem is a little different...


The total possible combinations of letters and envelopes is 3! = 6

Think of it this way, the first letter (L1) has 3 choices (E1, E2, E3). So say we match L1 to E1. The second letter will have only 2 choices (E2, E3) left since E1 is already taken by L1. Say we match L2 to E2. Then the final letter has only 1 choice left (E3) as the other 2 are already taken.

Lets list the possible options...

E1L1 E2L2 E3L3 --> Atleast 1 letter in proper envelope
E1L1 E2L3 E3L2 --> Atleast 1 letter in proper envelope
E1L2 E2L1 E3L3 --> Atleast 1 letter in proper envelope
E1L2 E2L3 E3L1
E1L3 E2L1 E3L2
E1L3 E2L2 E3L1 --> Atleast 1 letter in proper envelope

Thats a total of 6 possible combinations. Can anyone list any additional combinations thats not already listed above?

Also, as you can, 4/6 have atleast 1 letter in proper envelope

Hence probability is 2/3.

Hope this helps.

Cheers.
dtweah wrote:
maihuna wrote:Humm...your answer is incorrect, see yourself then where your fundamental principle of counting is putting you in trouble....
"An Envelope in a letter is different from an E's L's Combination." Why?

E1 E2 E3 L1 L2 L3

L1 can go in E1 , that is one possibility

L1 can go into E2, that is another possibility. What prevents this in the wording of the problem?

L1 can go into L3, a third possibility. What prevents this?
Attempt 1: 710, 92% (Q 42, 63%; V 44, 97%)
Attempt 2: Coming soon!
Join the discussion

by kaulnikhil » Fri Aug 14, 2009 12:49 pm
take letters A B and C
probability they wont fit in right envelope ..
A has two options ---- 2/3
B has 1 option ----1/2
C has 2 fit in the last one --1
multiply 2/3*1/2*1 = 1/3
therefore probability they fit = 1-1/3 = 2/3
Join the discussion